OpenStudy (anonymous):

4p − 2q = 16, 5p + 7q = 1

OpenStudy (anonymous):

in the first eqn 2p-q=16 in second 5p+7q=1 multiply 1sst by then add both eqn, you get 19p=56 then solve

OpenStudy (anonymous):

sry in 1st its =8

OpenStudy (anonymous):

2p-q=8 5p+7q=1 when eqn 1 multiplied by 7 and both are added, gives you 19p=56

OpenStudy (anonymous):

sry 19p=57

OpenStudy (anonymous):

solve the top for q by first subtracting 4 p -2q=-4p+16 divide by -2 q=2p-8 put this into the second equation 5p+7(2p-8)=1 multiply it out and add like terms 5p+14p- 56 = 1 19p-56=1 add 56 to both sides 19p-56+56=1+56 19p=57 divide by 19 to get p=3 then use this in either the first or second equation to find q (i chose the 2nd one) 5p+7q=1 5(3)+7q=1 15+7q=1 minus 15 both sides -15+15+7q=1-15 7q=-14 divide by 7 to get q=-2 check 5(3)+7(-2)=1 15-14=1 1=1 correct your answers are p=3 q=-2

OpenStudy (anonymous):

you get p=3 then substitute p in the first eqn you get q=2p-8 =6-8 =-2 therefore, (p,q)=(3,-2)

OpenStudy (anonymous):

thanks guys